i want to solve thisimultaneous equation 5x-5y=5,4x-3y=25
i took the first equation x=5+5y/5 and them
4(5-5y)/5-3y=25
i donīt understand the siplpification

Answer by Alan3354(69443)

i want to solve this simultaneous equation
It's a system of equations, 2 of them. You can't have one simultaneous anything.
----------------
5x-5y=5 --> x-y = 1 --> 3x-3y=3
4x-3y=25
3x-3y=3
-------- Subtract
1x = 22
x = 22
y = 21